# Upper Mid-Band Channel Measurements and Characterization at 6.75 GHz FR1(C) and 16.95 GHz FR3 in an Indoor Factory Scenario

### v1

@@ III-B RMS Delay Spread
Omni RMS $\lg_{\text{ASA}}$ $=\log_{10}$($\text{ASA}/1^{\circ}$) $\mu_{\text{lgASA}}^{\text{LOS}}$ 1.25 1.62 23.84 46.56 1.05 1.55 14.13 40.86
@@
$\mu_{\text{lgASA}}^{\text{NLOS}}$ 1.80 1.72 66.91 58.21 1.61 1.72 50.42 58.21
$\sigma_{\text{lgASA}}^{\text{NLOS}}$ 0.21 0.3 0.43 0.3
Omni RMS $\lg_{\text{ASD}}$ =$\log_{10}$($\text{ASD}/1^{\circ}$) $\mu_{\text{lgASD}}^{\text{LOS}}$ 1.43 1.56 36.01 39.01 1.34 1.56 22.88 39.01
@@
$\mu_{\text{lgASD}}^{\text{NLOS}}$ 1.64 1.57 46.65 38.90 1.49 1.57 34.72 38.90
$\sigma_{\text{lgASD}}^{\text{NLOS}}$ 0.21 0.2 0.32 0.2
@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles in azimuth and elevation planes. RMS AS is evaluated for a spatial lobe threshold of 10 dB below the strongest pointing direction, using procedures detailed in [22]. Table V summarizes the AS observed for both arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.84 at 6.75 GHz and 14.13 at 16.95 GHz. Wide ASAs of 66.91 and 50.42 are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP models a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [13]. ASAs are observed to decrease with frequency in both LOS and NLOS. Further, the ASDs are observed as 36.01 in LOS and 46.65 in NLOS at 6.75 GHz, while ASDs of 22.88 in LOS and 34.72 in NLOS are observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39 in LOS and 38.9 in NLOS across frequencies.
Point data for large-scale spatio-temporal parameters corresponding to each of the TX-RX location pairs is summarized in Table VI. The format and procedure follows the detailed approach shown for the indoor hotspot environment in [21].

### v2

@@ III-B RMS Delay Spread
Omni RMS $\lg_{\text{ASA}}$ $=\log_{10}$($\text{ASA}/1^{\circ}$) $\mu_{\text{lgASA}}^{\text{LOS}}$ 1.25 1.62 23.71 46.56 1.05 1.55 14.32 40.86
@@
$\mu_{\text{lgASA}}^{\text{NLOS}}$ 1.80 1.72 66.38 58.21 1.61 1.72 50.40 58.21
$\sigma_{\text{lgASA}}^{\text{NLOS}}$ 0.21 0.30 0.43 0.30
Omni RMS $\lg_{\text{ASD}}$ =$\log_{10}$($\text{ASD}/1^{\circ}$) $\mu_{\text{lgASD}}^{\text{LOS}}$ 1.43 1.56 35.89 39.02 1.34 1.56 22.91 39.02
@@
$\mu_{\text{lgASD}}^{\text{NLOS}}$ 1.64 1.57 45.93 38.90 1.49 1.57 34.77 38.90
$\sigma_{\text{lgASD}}^{\text{NLOS}}$ 0.21 0.20 0.32 0.20
@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles in azimuth and elevation planes. RMS AS is evaluated for a spatial lobe threshold of 10 dB below the strongest pointing direction, using procedures detailed in [22]. Table V summarizes the AS observed for both arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.71 at 6.75 GHz and 14.32 at 16.95 GHz. Wide ASAs of 66.38 and 50.40 are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP models a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [13]. ASAs are observed to decrease with frequency in both LOS and NLOS. Further, the ASDs are observed as 35.89 in LOS and 45.93 in NLOS at 6.75 GHz, while ASDs of 22.91 in LOS and 34.77 in NLOS are observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39 in LOS and 38.9 in NLOS across frequencies.
Point data for large-scale spatio-temporal parameters corresponding to each of the TX-RX location pairs is summarized in Table VI. The format and procedure follow the detailed approach shown for the indoor hotspot environment in [21].